# Online Navigation Planning for Long-term Autonomous Operation of Underwater Gliders

### v1

@@ Abstract
Underwater glider robots have become an indispensable tool for ocean sampling. Although stakeholders are calling for tools to manage increasingly large fleets of gliders, successful autonomous long-term deployments have thus far been scarce, which hints at a lack of suitable methodologies and systems. In this work, we formulate glider navigation planning as a stochastic shortest-path Markov Decision Process and propose a sample-based online planner based on Monte Carlo Tree Search. Samples are generated by a physics-informed simulator that captures uncertain execution of controls and ocean current forecasts while remaining computationally tractable. The simulator parameters are fitted using historical glider data. We integrate these methods into an autonomous command-and-control system for Slocum gliders that enables closed-loop replanning at each surfacing. The resulting system was validated in two field deployments in the North Sea totalling approximately 3 months and 1000 km of autonomous operation. Results demonstrate improved efficiency compared to straight-to-goal navigation and show the practicality of sample-based planning for long-term marine autonomy.
I Introduction
@@
Ocean sampling is the process of measuring indicators such as temperature, salinity, oxygen, and biomass. It is a crucial task with many applications in the sciences including biology, ecology, climatology, and meteorology. Findings based on these observations can inform policy around societal issues such as the management of the resources and ecosystems of the ocean. Ocean sampling is inherently challenging, since it involves measuring a turbulent fluid over a range of possible scales [1].
@@
Underwater gliders, a type of autonomous underwater vehicle (AUV), are central to addressing the ocean observation problem. In contrast to static networks of ocean floats [2, 3], gliders allow for the sampling of dynamic phenomena such as eddies or ocean fronts. Gliders are able to operate autonomously for months at a time and can be much more energy-efficient and less carbon-intensive than ship operations.
@@
Gliders now form an important part of the Global Ocean Observation System, and their activity has been increasing in the last decade [4]. It is widely accepted that their operation and management need to become more efficient in order for organisations to control increasingly large fleets of gliders [5, 6]. Standard operational procedures require input from human pilots upon the surfacing of the glider, which occurs every few hours. Pilots check weather conditions, the progress of the previous dive, and mission objectives to decide the next control instructions. This raises critical challenges as piloting supervision is not always available and the glider will continue to follow its most recent, possibly outdated, control instructions. Therefore, opportunities for optimising navigation planning are regularly lost.
@@
Figure 1: Slocum glider being deployed during a preliminary experiment. The proposed Glider Long-term Autonomy Engine (GLAE) system was used to pilot the glider autonomously in two long-term deployments totalling approximately $3$ months and $1000$ km.
@@
A promising direction is to leverage navigation planning algorithms through autonomous command-and-control systems. Such systems can plan online at every surfacing given the up-to-date surfacing location and environmental conditions. Despite the existence of a fairly large body of work on navigation planning for underwater gliders (reviewed in Section II), until now such methods have not seen wide adoption for autonomous long-term missions. We argue that this is primarily due to misalignment between problem formulations and operational realities, as well as an inability to account for noise in ocean current forecasts and the execution of control commands. Furthermore, in practical terms, while existing software enables command execution and scripted missions, it does not support online replanning towards long-term mission goals under environmental and actuation uncertainty.
@@
In this work, we formulate the glider navigation planning problem as a stochastic shortest-path Markov Decision Process in which dives correspond to discrete decision steps and transitions are impacted by uncertainty in the currents and glider motion. Our approach combines a computationally efficient physics-based simulator that generates possible post-dive states with an online sample-based Monte Carlo Tree Search planner. These components are integrated in the Glider Long-term Autonomy Engine (GLAE) system, which enables online, closed-loop replanning at every surfacing. We evaluate the approach in simulation and two long-term field trials in the North Sea totalling approximately $3$ months and $1000$ km of autonomous operation (a photograph from a preliminary deployment is shown in Figure 1). The results demonstrate improved efficiency relative to standard navigation. To our knowledge, our work is the most extensive fully autonomous field evaluation of online glider navigation planning to date.
@@
  • MDP formulation for glider navigation planning under uncertainty. We model multi-dive glider navigation planning as a stochastic shortest-path Markov Decision Process that explicitly accounts for uncertainty in ocean current forecasts and execution of controls.
@@
  • A fully integrated autonomous command-and-control system for Slocum gliders. We implement the proposed methods in the GLAE system, which is capable of controlling commercial Slocum gliders.
@@
  • Extensive validation. We evaluate the approach in two North Sea campaigns totalling over 33 months and $1000$ km of autonomous operation. Compared to standard navigation, in simulation we obtain gains of up to $9.88\%$ in dive duration and $16.51\%$ in transect length on average. We also achieve a statistically significant transect length reduction of $9.55\%$ in a field deployment.
II Prior Work on Glider Navigation Planning
There are two main categories of previous techniques for glider navigation planning: those based on optimal control and those relying on discrete search.
II-1 Optimal Control
@@
In the first (and most substantial) category, an optimal control approach for planning a trajectory between two waypoints was proposed in [7], where the problem is transformed into a nonlinear programming formulation and solved via numerical optimisation. An extension for handling time-varying currents was introduced in [8]. A path planning approach for estuarine environments (characterised by fast currents) that can optimise for both speed and energy usage was proposed in [9]. Ideas from fluid dynamic equations were combined with control theory in [10] to yield a general method for vehicles traveling through a variable flow field. Most recently, [11] proposed a method for trajectory planning for AUVs with active propulsion using high-fidelity ocean forecasts, applying it to an estuarine environment.
@@
An overall challenge for such path planning methods is that they assume deterministic motion and environments. As a result, they cannot account for or reason about noise in the forecasts and the movement of the glider, which can lead to compounding errors. Furthermore, while they can achieve impressive performance in minimising cost to destination, this is demonstrated in simulation and not in the field. It is therefore difficult to judge how well the proposed models align with reality.
It is also worth mentioning a line of work that considers the coordination of a fleet of several gliders for achieving effective sampling. We note, however, that the objective in these works is different from planning an efficient route to a destination. A control law that allows a formation of three gliders to maintain an equilateral triangle structure was formulated in [12], which can be used to determine gradients of properties such as temperature and salinity. A general approach for the control of a mobile sensor network such as a fleet of gliders was developed in [13]. The latter method was successfully applied in the ASAP trial [14], which deployed $6$ Slocum gliders over $24$ days covering a total distance of $3270$ km. However, this required $14$ interventions from human pilots to adjust the motion patterns based on weather forecasts.
II-2 Discrete Search
@@
In this second category of approaches, path planning using a variant of the A* algorithm was proposed in [15]. To render the algorithm applicable, it imposes a constant dive duration and is limited to handling 2D current fields. The work in [16] is perhaps closest to ours, given that it adopts a Markov Decision Process problem formulation, considers currents across multiple depth levels, and conducts a field trial. However, a highly limiting discretisation is applied to the state space, which allows for offline planning using a value iteration algorithm. Furthermore, the field experiment is limited to a single dive.
@@
In summary, our work is among the first to couple uncertainty-aware online navigation planning with extended real-world deployments of ocean gliders. The multi-month campaigns presented here provide one of the most extensive demonstrations to date of fully autonomous glider navigation without regular human-in-the-loop interventions.
III Background
Figure 2: An illustration of the key inputs and outputs for the glider dive simulator, which is used to generate stochastic MDP transitions.
@@ III Background
III-A Underwater Gliders
Underwater gliders [1] are AUVs that move horizontally on wings and vertically by controlling buoyancy. Descent and ascent are achieved via a piston or an artificial external bladder, changing the volume of the AUV but keeping mass constant. Movement is performed through dives consisting of several yos, forming a sawtooth pattern. At the start of a yo, the volume is decreased and the vehicle descends towards the ocean floor; at the bottom of each yo, the volume is increased and the vehicle ascends. The angle of the dive is controlled by shifting mass (i.e., the battery) along the internal tubing. The efficient passive propulsion of gliders renders them capable of operating for months at a time.
Gliders can be equipped with a large array of sensors for ocean sampling. Commonly used sensors include temperature, salinity, and pressure. They also feature a GPS and satellite antenna, allowing them to transmit sampled scientific data and receive control instructions. The instructions are comprised of the waypoints that the glider should aim to reach and the control parameters. Notable control parameters include the yo floor $z_{\text{bottom}}$, ceiling $z_{\text{top}}$, number of yos $n_{\text{yos}}$, dive pitch $\theta_{\text{dive}}$, climb pitch $\theta_{\text{climb}}$, and buoyancy change $\chi$. The control parameters determine the behaviour of the lower-level controller on the glider during the dive. A notable challenge is that the transfer of control instructions, which happen over satellite, are not fully reliable. Disruptions can occur in periods with adverse weather, and it is common for several transfer failures to happen in a row. The glider will use the latest set of instructions that transferred successfully, which may be significantly out of date.
Standard piloting relies on manually changing the control instructions. In between surfacings, pilots check several data sources indicating glider state and environmental factors such as wind and waves in order to prepare the next set of control instructions. Given this process occurs several times a day for months, it is not feasible for pilots to monitor every surfacing. Therefore, pilots generally set the closest waypoint such that several dives will be required to achieve it. The glider recomputes the heading towards this waypoint upon every surfacing, a mechanism we refer to as straight-to-goal (STG). By diving straight towards the waypoint, opportunities are regularly lost to optimise considerations such as mission duration (e.g., setting a heading that exploits favorable currents). This motivates our investigation of online navigation planning techniques, which can adapt to changes in glider state and environmental factors after every surfacing.
III-B MDP Preliminaries
Markov Decision Processes (MDPs) are a formalisation of decision-making [17, 18]. We define an MDP as the tuple $(\mathcal{S,A,T,C,G})$, where $\mathcal{S}$ is the state space (all configurations the agent may find itself in), $\mathcal{A}$ is the action space defining valid controls, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ denotes the transition function, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,\infty)$ is a cost function. Also let $\mathcal{G}\subseteq\mathcal{S}$ denote a set of absorbing goal states s.t. $\forall s_{g}\in\mathcal{G}\land s^{\prime}\notin\mathcal{G}$ we have $\mathcal{T}(s_{g},a,s_{g})=1$, $\mathcal{T}(s_{g},a,s^{\prime})=0$, and $\mathcal{C}(s_{g},a,s_{g})=0$. This class of models is known as Stochastic Shortest-Path (SSP) MDPs.
A deterministic, stationary, Markovian policy is a mapping $\pi:\mathcal{S}\to\mathcal{A}$. The value function of such a policy for a state $s$ is defined as $V^{\pi}(s)=\mathbb{E}[\sum_{t^{\prime}=0}^{\infty}{C_{t^{\prime}+t}^{\pi_{s}}}]$, where $C_{t^{\prime}+t}^{\pi_{s}}$ is a random variable denoting the cost incurred as a result of executing policy $\pi$ starting in state $s$ for all $t$. The optimal value function $V^{*}$ satisfies $V^{*}(s)=\min_{\pi}V^{\pi}(s)$, i.e., it specifies the minimal cost attainable in expectation at each state. We seek to find the corresponding optimal policy $\pi^{*}$ which, for the class of SSP MDPs, is deterministic and Markovian (i.e., independent of history given the current state) [17]. Finding or approximating $\pi^{*}$ can be achieved through various MDP solution methods.
III-C MCTS for Online Planning
For many practical problems, the size of the MDP is very large [19, 17], prohibiting us from solving it exhaustively and optimally in practice. Thus, we often have to resort to solving the MDP approximately from the current state. In this work, we leverage Monte Carlo Tree Search (MCTS), an online search algorithm for determining the best action to apply at the current state of an MDP. As a full review of MCTS is out of scope, we refer the interested reader to [20] for more details.
Upper Confidence Bounds for Trees (UCT) [21] is a widely adopted MCTS algorithm. It treats action selection at each node as an independent multi-armed bandit problem, applying the principles behind the Upper Confidence Bounds (UCB) [22] algorithm to the tree search setting. UCT balances exploration of actions that have not yet been trialled with exploitation of actions that have lead to high returns. Following the principle of optimism in the face of uncertainty, the node corresponding to the action with the highest upper limit of a confidence interval on returns is chosen to be visited next.
When a node with unexplored actions is visited, MCTS creates a new node and runs a trial to estimate its value. The result of the trial is backpropagated towards the root of the search tree, updating the node statistics. This process is repeated until the computational budget, expressed in terms of number of trials or wall clock time, is exhausted. The depth of the search tree and quality of the MCTS solution strongly depend on the number of trials that can be ran at each planning step. When the computational budget is exhausted, the algorithm selects the child node of the root with the maximum visit count or best estimated value. The corresponding action will be executed in the environment.
Many variants and enhancements on the basic MCTS framework have been proposed. Notably, double progressive widening is suitable for scenarios characterised by stochastic transitions and continuous state spaces [23, 24]. It is able to achieve a suitable compromise between bias (induced by the first expanded nodes) and variance (in the result of trials for each state).
Parallelisation techniques have also been proposed in order to take advantage of multi-core architectures despite the sequential nature of MCTS. Notably, the root-parallel variant [25, 26] creates a number of search trees that progress independently from the given root state. The final results are aggregated (e.g., by majority voting) to obtain a single action.
IV Methods
IV-A Problem Statement

### v2

@@ Abstract
Underwater glider robots have become indispensable for ocean sampling, yet fully autonomous long-term operation remains rare in practice. Although stakeholders are calling for tools to manage increasingly large fleets of gliders, existing methods have seen limited adoption due to their inability to account for environmental uncertainty and operational constraints. In this work, we demonstrate that uncertainty-aware online navigation planning can be deployed in real-world glider missions at scale. We formulate the problem as a stochastic shortest-path Markov Decision Process and propose a sample-based online planner based on Monte Carlo Tree Search. Samples are generated by a physics-informed simulator calibrated on real-world glider data that captures uncertain execution of controls and ocean current forecasts while remaining computationally tractable. Our methodology is integrated into an autonomous system for Slocum gliders that performs closed-loop replanning at each surfacing. The system was validated in two North Sea deployments totalling approximately 3 months and 1000 km, representing the longest fully autonomous glider campaigns in the literature to date. Results demonstrate improvements of up to 9.88% in dive duration and 16.51% in path length compared to standard straight-to-goal navigation, including a statistically significant path length reduction of $9.55\%$ in a field deployment.
{IEEEkeywords}
ocean gliders, navigation planning, Monte Carlo Tree Search, long-term autonomy
1 INTRODUCTION
[…]
@@ 1 INTRODUCTION
Ocean sampling is the process of measuring indicators such as temperature, salinity, oxygen, and biomass. It is a crucial task with many applications in the sciences including biology, ecology, oceanography, climatology, and meteorology. Findings based on these observations can inform policy around societal issues such as the management of marine resources and ecosystems. Ocean sampling is inherently challenging, since it involves measuring a dynamic fluid over a range of temporal and spatial scales [1].
@@
Underwater gliders, a type of autonomous underwater vehicle (AUV), are central to addressing the ocean observation problem. In contrast to static networks of hydrographic moorings or drifting networks of floats [2, 3], gliders allow for the sampling of dynamic phenomena such as eddies or ocean fronts. Gliders are able to operate autonomously for months at a time and can be more energy-efficient and less carbon-intensive than ship operations [4].
@@
Gliders form an important part of the Global Ocean Observation System, and their activity has been increasing in the last decade [5]. It is widely accepted that their operation and management need to become more efficient in order for organisations to control increasingly large fleets [6, 7]. Standard operational procedures require input from human pilots upon the surfacing of the glider, which occurs every few hours. Pilots check environmental conditions, the progress of the previous dive, and mission objectives to decide the next control instructions. This raises challenges as piloting supervision is not always available, leading the glider to follow its most recent, possibly outdated, control instructions. Therefore, opportunities for optimising navigation are regularly lost.
@@
Figure 1: Slocum glider in operation during the eSWEETS3 field deployment. The proposed Glider Autonomy Long-term Planning Engine system was used to pilot gliders autonomously in two long-term deployments totalling approximately 33 months and $1000$ km.
@@
A promising direction to optimise navigation autonomously is to leverage navigation planning algorithms, which can plan online at every surfacing given an up-to-date surfacing location and environmental conditions. Despite the existence of a fairly large body of work on navigation planning for underwater gliders (reviewed in Section 3), until now such methods have not been adopted for autonomous long-term missions. We argue that this is primarily due to the inability of such methods to account for uncertainty in ocean current forecasts and vehicle dynamics.
@@
In this work, we formulate the glider navigation planning problem as a Markov Decision Process (MDP) in which dives correspond to actions, and in which transitions capture uncertainty in the ocean currents and glider motion. Our approach combines a computationally efficient physics-based simulator that generates possible post-dive states with an online sample-based Monte Carlo Tree Search planner. These components are integrated in the Glider Autonomy Long-term Planning Engine (GALE) system, which enables online, closed-loop replanning at every surfacing.
GALE was used to conduct two operational deployments in the North Sea towards refining weather forecasts and understanding the environmental impact of offshore wind farms. Our field evaluation, which exceeded $3$ months and $1000$ km in total, represents the most extensive fully autonomous deployment of online glider navigation planning to date. Results from the field evaluation and simulations demonstrate improved efficiency relative to the standard navigation approach.
@@
  • Markov Decision Process formulation for glider navigation planning under uncertainty. We model multi-dive glider navigation planning as a stochastic shortest-path MDP that explicitly accounts for uncertain ocean current forecasts and glider motion.
@@
  • The design of an autonomous online planning system for Slocum gliders. We contribute the design of the GALE system, which is underpinned by the proposed methods. It is capable of controlling commercial Slocum gliders [8, 9].
@@
  • Large-scale validation in operational conditions. We evaluate the approach in two North Sea campaigns totalling over $3$ months and $1000$ km of autonomous operation, representing the most extensive real-world validation of online autonomous glider navigation planning to date. Compared to standard navigation, in simulation we obtain decreases of up to $9.88\%$ in dive duration and $16.51\%$ in path length on average. We also achieve a statistically significant path length reduction of $9.55\%$ in a field deployment.
2 UNDERWATER GLIDERS
Figure 2: An illustration of basic glider operation and the relevant control parameters for Slocum gliders. These include the yo floor $z_{\text{bottom}}$, ceiling $z_{\text{top}}$, climb pitch $\vartheta_{\text{climb}}$, dive pitch $\vartheta_{\text{dive}}$, buoyancy change $\Delta_{\text{buoy}}$, and number of yos $n_{\text{yos}}$ (33 yos are completed in this dive).
An underwater glider [1] is a type of AUV that moves through the ocean by changing its buoyancy to glide up and down through the water column. Descent and ascent are achieved via a piston or an external bladder, changing the volume of the AUV but keeping mass constant. Movement is performed through dives consisting of several yos, forming a sawtooth pattern as shown in Figure 2. At the start of a yo, the volume is decreased and the vehicle descends towards the ocean floor; at the bottom of a yo, the volume is increased and the vehicle ascends. The glider’s wings convert some of the vertical motion into forwards propulsion. Controlling the angle of the dive is achieved by shifting internal masses (typically the battery). Gliders are capable of operating for months at a time given their efficient passive propulsion.
Gliders are equipped with sensors for measuring indicators such as temperature, salinity, and pressure. They also feature a GPS and satellite antenna, allowing them to transmit sampled scientific data and receive control instructions. The instructions are comprised of two elements: the waypoints that the glider should aim to reach, and the control parameters that determine the behaviour of the lower-level controller on the glider during the dive. The control parameters are illustrated in Figure 2.
The online planning loop is illustrated in Figure 3. Control instructions are transferred over satellite while the glider is floating at the surface. The time spent floating is dictated by the amount of scientific data to be transferred and is on the order of 5-10 minutes. This implicitly defines a time budget for determining the next control instructions. An additional challenge is that transfers of control instructions are not fully reliable. Disruptions can occur during periods with adverse weather, and it is common for several transfer failures to happen in a row. The glider will use the latest set of instructions that transferred successfully, which may be significantly out of date.
Standard piloting relies on manually changing the control instructions. In between surfacings, pilots check several data sources indicating glider state and environmental factors such as wind, waves, and ocean currents in order to prepare the next set of control instructions. Given the glider surfaces several times a day for months, it is not feasible for pilots to monitor every surfacing. Despite supervision only being intermittent, estimates from recent deployments suggest that manual piloting accounts for approximately 18% of total campaign costs, which organisations are keen on reducing.
Pilots generally set the closest waypoint such that several dives will be required to achieve it. The glider recomputes the heading towards this waypoint upon every surfacing, a mechanism we refer to as straight-to-goal (STG). By diving straight towards the waypoint, opportunities are regularly lost to optimise considerations such as mission duration (e.g., by setting a heading that exploits favorable currents). This motivates our investigation of online navigation planning techniques, which can adapt to changes in glider state and environmental factors after every surfacing.
Figure 3: The online planning loop of the GALE system and key considerations. Upload Science Data and Plan Next Dive occur simultaneously while the glider is at the surface.
3 RELATED WORK
Glider navigation planning has been addressed under several distinct problem formulations. Works based on optimal control focus on devising an optimal path under known flow fields. A second category of approaches relies on discrete Markov Decision Process formulations that are solved via search or learning. Lastly, another category of works focuses on optimising coordination and coverage for fleets of gliders, rather than planning an efficient route to a destination.
3.A Optimal Control
@@
Planning a trajectory between two waypoints was studied in [10], where the problem is transformed into a nonlinear programming formulation and solved via numerical optimisation. An extension for handling time-varying currents was introduced in [11]. A path planning approach for estuarine environments (characterised by strong tidal currents) that can optimise for both speed and energy usage was proposed in [12]. Ideas from fluid dynamic equations were combined with control theory in [13] to yield a general method for vehicles traveling through a variable flow field. Most recently, [14] proposed a method for trajectory planning for AUVs with active propulsion using high-resolution ocean forecasts, applying it to an estuarine environment.
@@
A key limitation of such path planning methods is that they assume deterministic motion and environments. As a result, they do not explicitly model uncertainty in ocean current forecasts and the movement of the glider, which can lead to compounding errors over multi-dive horizons. Furthermore, while they can achieve impressive performance in minimising cost to destination, their evaluation is typically limited to simulation. This stands in contrast to our work, which explicitly models uncertainty and evaluates performance in multi-month field campaigns.
3.B Search and Markov Decision Process Formulations
@@
Path planning using a variant of the A* algorithm was proposed in [15]. To make the execution of A* tractable, the method assumes deterministic motion under a single realisation of a 2D current field and imposes a constant dive duration. The work in [16] adopts a Markov Decision Process problem formulation, considers currents across multiple depth levels, and conducts a field trial. However, this approach uses a highly limiting discretisation of the state space, which allows for offline planning using a value iteration algorithm. Uncertainty is incorporated through a Gaussian model over the summed vehicle and current velocities, yielding transition probabilities for each of the 8 surrounding cells. The approach assumes a fixed underlying current field and models uncertainty in a simplified manner, without accounting for forecast uncertainty or state-dependent execution noise. Furthermore, the field experiment is limited to two round-trip traversals between a pair of waypoints separated by $3.5$ km.
Very recent work [17] has applied deep reinforcement learning (RL) to glider path planning. The proposed approach uses ocean current forecasts, which are paired with a simplified glider dynamics model to yield a simulation environment. The path planning policy is trained using a Soft Actor-Critic architecture for optimising a complex multi-objective reward. While improvements are demonstrated in simulation over other deep RL variants, important limitations remain: the planning occurs entirely offline and is not adjusted based on the live surfacing location of the vehicle; the method does not incorporate forecast or execution uncertainty; and the field trial is limited to a single trajectory.
3.C Fleet Coordination
A control law allowing a formation of three gliders to maintain an equilateral triangle structure was formulated in [18]. It can be used to determine 3-dimensional gradients of ocean properties such as temperature and salinity. A general approach for controlling a mobile sensor network such as a glider fleet was developed in [19]. The latter method was successfully applied in the Adaptive Sampling and Prediction (ASAP) trial [20], which deployed $6$ Slocum gliders over $24$ days covering a total distance of $3270$ km. However, this required $14$ interventions from human pilots to adjust the motion patterns based on ocean forecasts. The work [21] considered the problem of planning paths for a fleet of gliders such as to achieve complete coverage of the sampling area. Ant colony optimisation was used for altering preset paths in order to avoid obstacles. In simulation, the method yielded shorter paths with less consumed energy. We note that these approaches focus on the mission-level objectives, and are complementary to our work, which addresses the lower-level problem of navigation planning
3.D Summary
@@
Overall, our work is the first to couple uncertainty-aware online navigation planning with extended real-world deployments of ocean gliders. To our knowledge, the multi-month campaigns presented here are the most extensive demonstration to date of fully autonomous glider navigation without regular human-in-the-loop interventions.
Figure 4: An illustration of the key inputs and outputs of the glider dive simulator. The glider icon represents the dive initial location, the arrow indicates its direction, and background arrows correspond to ocean current forecasts over several depth levels. The grey circles are the outputs of the simulator indicating possible surfacing locations, i.e., stochastic MDP transitions. The red triangle is the true surfacing location.
@@
4 METHODS
4.A Problem Statement